OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

OpenStudy (anonymous):

Find slope m = (y2-y1)/(x2-x1) (x1,y1), (x2,y2) are given points then use point slope formula (y-y1)= m(x-x1)

OpenStudy (michele_laino):

@kaylamarie081 Please you have to substitute your coordiates in all of your equation, which you have wrote. When you get an identity, you have found the right equation

OpenStudy (michele_laino):

@kaylamarie081 please, you have to substitute both your points in each of your equation

OpenStudy (anonymous):

sangya21 is right, but then you might have to convert your answer to the format the answer choices are in. Do you need help with that?

OpenStudy (anonymous):

Yeah I really sick at math and no clue what I'm doing

OpenStudy (anonymous):

Suck*

OpenStudy (anonymous):

(x1,y1) = (-1, 3) (x2,y2) = (1,11) m = (y2-y1)/(x2-x1) m = (11-3)/(1-(-1)) m = 8/2 = 4 then use point slope formula (y-y1)= m(x-x1) (y-3)= 4(x-(-1)) y - 3 = 4x +4 y = 4x +7

OpenStudy (anonymous):

Is my answer B?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Its D 4x - y = -7 4x+7 = y